# ArtWhisperer: A Dataset for Characterizing Human-AI Interactions in Artistic Creations

### v1

@@ 3.2 People submit similar prompts throughout their interaction
While this is not a surprising result (that users often do not make significant changes to their prompt), but it is an important result to understand how typical users interact with AI models. Moreover, it suggests that user initialization (i.e., the first prompt they submit) is critical.
@@ 3.3 People have similar prompt styles across images
For each user and each target image, we calculate the average prompt embeddings and the average prompt embedding submitted by the given user. We take the difference of these two embeddings as a user specific average prompt, that represents the style of the user. We then compute the standard deviation these user specific embeddings across target images, as a measure of how much a user varies in prompting style across images.
In the right of Figure 5, we plot the distribution of the standard deviations for users (blue) and for permuted users (orange), where the permuted users are generated by sampling user specific average prompts and assigning them to the simulated players, allowing us to test whether real players have more similarity than random across target images.
@@
Figure 4: Diverse, high-scoring prompt submissions from players. 4 examples of high-scoring prompts from different users. Target image shown in green column. Plots show image (left) and text (right) embeddings of the displayed images (target in orange; user submissions in green; first submission by corresponding user in red) using the Umap [15] projection.
Figure 5: Left, Diversity of prompts does not diminish as users improve their score with better quality prompts. Center, Individual users have similar prompts for a given image. Right, Individual users have different styles of prompting.
@@ 4 Model Steerability
Model steerability refers to the ability of a user to steer a model towards a desired outcome. There is no current consensus on how to measure AI steerability. A common approach is to simply measure performance of a model on standardized dataset evaluations [10, 17]. While this can enable comparisons between tasks and models, this approach does not allow for the feedback loop present when humans interact with a model. Steerability can also be measured qualitatively based on user assessment of their experience interacting with the AI [5]. We create a simple yet informative measure of model steerability. We then analyze this measure across different subgroups of images and across two different Stable Diffusion models–SDv2.1 and the older SDv1.5 [24].
@@ 4.2 Analysis
In Figure 6, we plot the steerability score across image groups. Error bars show the standard error. For examples of steerability scores for individual images, see the Supplement. We find that images containing famous people or landmarks, real images (not AI generated), contain cities, or contain nature are the most steerable. AI-generated images, fantasy images, and images of human art are the least steerable. There are a few possible explanations. The model we are assessing here, SDv2.1, as well as its text encoder OpenCLIP, are trained on subsets of LAION5B [27]. The contents of LAION5B are predominantly real world images, indicating why these images may be more steerable (i.e., text describing these types of images may have a better encoding). Moreover, the prompts for AI-generated images and fantasy images generally include specific internet artists and/or art styles which may not be known to most users making achieving the desired target image more difficult. Another potential reason is the distribution of images chosen for each category. Clearly, there are “easier” and “more difficult” images in each category; part of the reason for smaller stopping time may be the sample of images chosen rather than the actual image category.
We also compare steerability across the two models: SDv2.1 and SDv1.5. Across most image categories, we observe a similar steerability. Images of nature, sci-fi or space, and real images have the largest differences in steerability between the two models; SDv2.1 is more steerable in all three cases. This suggests that SDv2.1 may be more steerable for natural images as well as sci-fi images, and is similarly steerable for other kinds of images including AI-generated artwork. One explanation may be that most of our users were not aware of certain prompting strategies that help models generate more aesthetic images or certain art styles; it is possible that for experienced users, AI art images may be more steerable, and differences between models may be magnified if, for example, a user is experienced working with one particular model. More discussion is provided in the Appendix.
@@ 4.3 Justification for automated score
One limitation in our steerability metric comes from the method of scoring user-submitted prompts. Ideally, we would like to assess steerability based on a user’s personal preferences. As mentioned in Section 2.2, the scores and human ratings have a positive correlation. Here, we use the human ratings instead of our score function to assess steerability. We compute the steerability score across both models and across image groups. Generally, the steerability scores change little. In all but two cases (SDv2.1 on sci-fi and space images; SDv1.5 on nature images), the human rating-based steerability score remains within a 95% confidence interval of the score-based steerability score. While our score function may not perfectly capture human preferences, the steerability score we generate appears to be robust to these issues. Further discussion is included in the Appendix.
Figure 6: Steerability across image groups (smaller indicates increased steerability). Bars show average expected stopping time across images in the image group; error bars show standard error of the mean.
@@ 5 Discussion
As demonstrated in our initial analysis, the ArtWhisperer dataset can provide insights into user prompting strategies and enables us to assess model steerability for individual tasks and groups of tasks. What makes our dataset particularly useful is the controlled interactive environment, where users work toward a fixed goal, that we capture data in. We expect the ArtWhisperer dataset will be useful beyond the initial analyses presented here. For example, some future work may involve fine-tuning models to better accommodate users based on their prompt trajectories or guide users towards better prompting strategies. There is also deeper analysis possible for understanding user prompting strategies.
@@ Appendix A Appendix
A.1 Information on Wikipedia pages scraped
@@
A.2 Scoring function details
In Section 2.2, we defined the scoring function, $score(x_{i},t_{i})$ as
$score(x_{i},t_{i})=\max(0,\min(100,\alpha\cdot\frac{\langle CLIP(x_{i}),CLIP(t_{i})\rangle}{||CLIP(x_{i})||_{2}\cdot||CLIP(t_{i})||_{2}}+\beta)).$
@@ A.2 Scoring function details
Since our score range is $[0,100]$ but cosine similarity has a range of $[-1,1]$, we scale the cosine similarity by $100$, resulting in
@@
We also add a “score adjustment” term that attempts to normalize image difficulty. In particular, for each target image, we compute the un-clipped score for the target image $t_{i}$ and the image generated using the target prompt with, $x_{i}$:
$unclipped\_score(x_{i},t_{i})=\alpha^{\prime\prime}\cdot\frac{\langle CLIP(x_{i}),CLIP(t_{i})\rangle}{||CLIP(x_{i})||_{2}\cdot||CLIP(t_{i})||_{2}}+\beta^{\prime\prime}.$
This score assess the score a user would receive if they exactly entered the target prompt. We fix this value as $100$ (i.e., a perfect score prior to clipping), and set the score adjustment parameter, $c_{i}$, to appropriately normalize this score. In particular,
$c_{i}=\frac{100}{unclipped\_score(x_{i},t_{i})},$
@@
$\alpha_{i}$ $=-150.3\cdot c_{i}$
$\beta_{i}$ $=179.1\cdot c_{i}$
We use the target specific parameters, $\alpha_{i}$, $\beta_{i}$, for our scoring function parameters, $\alpha$, $\beta$, (so each target image may have slightly different parameters).
A.3 Additional information on running the game
@@
A.4 Additional example images
We provide additional examples of image trajectories and diverse images in Figures 8 and 9.
@@ A.4 Additional example images
Figure 9: More examples of diverse high-scoring prompts, as in Figure 4.
A.5 Algorithm for steerability
@@
Input: Set of Images
Output: Steerability estimate
Procedure EstimateStoppingTime(targetImages)
       Initialize array $steerability$ to track stopping times;
       Initialize $scoreBins\leftarrow[[0,20],[21,40],[41,60],[61,80],[81,100]]$;
       Initialize $\epsilon\leftarrow 1$;
[…]
             ${markov}_{i}$leftarrow EstimateMarkovChain(${image}_{i}$, $scoreBins$, $\epsilon$);
             $steerability[i]$leftarrow RunMonteCarloEstimation(${markov}_{i}$);
       end for
      return $\mathbf{E}[steerability]$
Procedure EstimateMarkovChain(targetImage, bins, $\epsilon$)
       Initialize bin pair count as $counts[(\text{bin}_{i},\text{bin}_{j})]\leftarrow\epsilon$, using transitions from a dummy node to model the first prompt submitted;
[…]
                   Convert $\text{score}_{i,j}$ to bin number, $\text{bin}_{i,j}$;
                   Increment $counts[(\text{bin}_{i-1,j},\text{bin}_{i,j})]$ by $1$;
             end for
       end for
      Normalize empirical transition counts to find empirical node transition probabilities;
       Define $markov_{\text{targetImage}}$ using the node transition probabilities;
       return $markov_{\text{targetImage}}$
Procedure RunMonteCarloEstimation(markovChain)
       Run Monte Carlo simulation to estimate time to reach the last bin for markovChain, starting from the dummy (initial) node;
       return Estimated stopping time;
Algorithm 1 Algorithm for estimating image steerability.
A.6 Steerability across models

### v4

@@ 3.2 People submit similar prompts throughout their interaction
Second, we find that people do not explore a wide range of prompt designs even when their initial prompts are not performing well. This suggests the lack of convergence in prompt design is inherent to the users’ preference to describe an image. It also suggests that user initialization (i.e., the first submitted prompt) is critical, and that online personalization may be possible (to adapt to the user’s writing style) given the more stable nature of an individual’s prompt design for a given target image.
@@ 3.3 People have similar prompt styles across images
Finally, we find that people have a measurable writing style or prompt design that appears across images; however, this prompt design is predominantly informed by the target image rather than any consistent style. This implies that online personalization across different target images may require a large number of images to be effective (across a small set of individual target images, the prompt design and writing style may still change significantly). This is in contrast to our earlier finding that suggests personalization for a given image may be possible given the relative stability of a user’s prompt design for a given target image.
@@
Figure 4: Diverse, high-scoring prompt submissions from different users. Target image in rightmost column.
Figure 5: Difference of distance from the first prompt to ground truth and distance from the last (best) prompt to ground truth for CLIP text (blue) and CLIP image embeddings (orange).
Figure 6: Left: Users submit diverse (across users) prompts, both at beginning and end of interaction. Center: Individual users do not submit diverse prompts. Right: Users have different prompting styles.
@@ 4 Model Steerability
Model steerability refers to the ability of a user to steer a model towards a desired outcome. Measuring steerability has great utility, as it enables quantitative tracking of the human usability of generative models. However, there is no current consensus on how to measure AI steerability. A common approach is to simply measure performance of a model on standardized dataset evaluations [17, 28]. While this can enable comparisons between tasks and models, this approach does not allow for the feedback loop present when humans interact with a model. Steerability can also be measured qualitatively based on user assessment of their experience interacting with the AI [9].
Here, we propose a simple yet informative measure to assess model steerability. This metric is general across model types and data modalities, and we are able to use it here to compare across image categories and models. We then analyze this measure across different subgroups of images and across two different Stable Diffusion models: SDv2.1 and the older SDv1.5 [37].
@@ 4.2 Analysis
In Figure 7, we plot the steerability score across image groups. Error bars show the standard error. For examples of steerability scores for individual images, see Appendix A.12. We find that images containing famous people or landmarks, real images (not AI generated), contain cities, or contain nature are the most steerable. AI-generated images, fantasy images, and images of human art are the least steerable. There are a few possible explanations. The model we are assessing here, SDv2.1, as well as its text encoder OpenCLIP, are trained on subsets of LAION5B [40]. The contents of LAION5B are predominantly real world images, indicating why these images may be more steerable (i.e., text describing these types of images may have a better encoding). Moreover, the prompts for AI-generated images and fantasy images generally include specific internet artists and/or art styles which may not be known to most users making achieving the desired target image more difficult. Another potential reason is the distribution of images chosen for each category. Clearly, there are “easier” and “more difficult” images in each category; part of the reason for smaller stopping time may be the sample of images chosen rather than the actual image category.
Using the ArtWhisperer-Validation data, we also compare steerability across two models: SDv2.1 and SDv1.5. Across most image categories, we observe a similar steerability. Images of nature, sci-fi or space, and real images have the largest differences in steerability between the two models; SDv2.1 is more steerable in all three cases. This suggests that SDv2.1 may be more steerable for natural images as well as sci-fi images, and is similarly steerable for other kinds of images including AI-generated artwork. One explanation may be that most of our users were not aware of certain prompting strategies that help models generate more aesthetic images or certain art styles; it is possible that for experienced users, AI art images may be more steerable, and differences between models may be magnified if, for example, a user is experienced working with one particular model. More discussion is provided in Appendix A.11.
@@ 4.3 Justification for automated score
One limitation of our steerability metric comes from the method of scoring user-submitted prompts. Ideally, we would assess steerability with a user’s personal preferences. As mentioned in Section 2.2, the scores and human ratings have a positive correlation. Here, we use the human ratings from ArtWhisperer-Validation instead of our score function to assess steerability. We compute the steerability score across both models and across image groups. Generally, the steerability scores change little. In all but two cases (SDv2.1 on sci-fi and space images; SDv1.5 on nature images), the human rating-based steerability score remains within a 95% confidence interval of the score-based steerability score. While our score function may not perfectly capture human preferences, the steerability score we generate appears to be robust to these issues. Further discussion is included in Appendix A.14.
Figure 7: Steerability across image groups for human users and GPT-4 (smaller indicates increased steerability). Bars show average expected stopping time across images in the image group; error bars show standard error.
5 Vision-Language Model Evaluation
We also use the ArtWhisperer dataset to evaluate vision-language models (VLMs) for their ability to incorporate feedback. In particular, we have two VLMs, GPT-4 and Gemini [28, 43], play the ArtWhisperer game across all target images. In this context, the models are interacting with a “tool” – an SD model. However, we can also consider this SD model as a proxy for a user whose preferences the model must adapt to over time.
A system prompt is crafted to inform the model about the game. A starting prompt is used to query the model for an initial prompt. The generated prompt is evaluated using the ArtWhisperer game pipeline – an image is generated and then scored. In the “Feedback” mode, this generated image and the ArtWhisperer score is then fed back to the VLM with a request for an updated prompt. This process is repeated until the model attains a perfect score or 20 attempts have been made. In “No feedback” mode, the VLM is not given any feedback and is just queried repeatedly without any conversation history given. More details on the prompts used are included in Appendix A.16.
In Figure 8, we compare the average score (across target images) trajectory of the VLM models over time. Here, the score is the ArtWhisperer game score, where larger values indicate a better prompt. For each of GPT-4 and Gemini, we have two trajectories. “GPT-4” and “Gemini” represent the prompting methodology described above and further detailed in Appendix A.16. The “No feedback” plots represent the average score across queries as well, as each query is independent; this is a baseline measure of the VLM’s ability to prompt Stable Diffusion given an image with no feedback.
When given feedback, GPT-4 consistently improves as it receives more feedback, with minimal improvement after about 15 rounds of feedback. This indicates the model is able to incorporate feedback well. In contrast, Gemini does not improve with feedback. This is expected – the Gemini model we evaluated on was not trained for multi-turn conversations (an API for such a model was not released at the time of writing).
In Figure 7, we compare steerability of the Stable Diffusion model with respect to GPT-4 and humans. Across all tasks, Stable Diffusion is at least as steerable with respect to GPT-4 as it is with humans. This result holds even when we normalize for initial score (i.e., to reduce the effect of baseline prompt writing ability and just examine ability to use feedback), indicating that GPT-4 is better at adapting to feedback compared to human users. This additional plot and more discussion is included in Appendix A.16.
Figure 8: Averaged trajectory of GPT-4 and Gemini. “No feedback” does not use any feedback from previously generated images.
@@ 6 Discussion
As demonstrated in our analysis, the ArtWhisperer and ArtWhisperer-Validation datasets can provide insights into user prompting strategies and enables us to assess model steerability for individual tasks and groups of tasks. What makes our dataset particularly useful is the controlled interactive environment, where users work toward a fixed goal, that we capture data in.
One of the most exciting use cases we see for our dataset is to create synthetic humans for prompt generation. For example, similar to the method described in Promptist [13], we imagine fine-tuning a large language model with our dataset to generate prompt trajectories (i.e., rather than an optimized prompt) using similar exploration strategies as a human prompter. These synthetic prompters could be based on multimodal models like OpenFlamingo [2] or text-only models and use score-feedback to condition the trajectory generation. As an initial proof-of-concept, we fine-tuned a MT0-large model [26] model on our dataset and found the fine-tuned model can indeed behave similarly to human users (see Appendix A.15). It is also feasible that a VLM like the ones assessed in Section 5 could be used as synthetic humans through few-shot prompting, obviating the need for fine-tuning. These synthetic prompters have several uses:
  1. 1.
    Automating measurement of text-to-image model steerability by using synthetic users in place of real human prompters. While we believe our proposed steerability metric is effective, its main limitation currently is the requirement for human annotations.
  2. 2.
    Incorporating steerability in the objective function for text-to-image models. By representing steerability as a function of synthetic users, it becomes possible to explicitly optimize a model for steerability.
  3. 3.
    Generating human readable image captions that are compatible with a Stable Diffusion model by using the synthetic prompter to optimize the token representation of the prompt. This is related to [54], where the authors use models to revise prompts.
Additionally, our dataset can be used for further analysis on human prompting strategies beyond what we discussed in the paper. For example, one question we only touched upon is whether we can compare human prompters to automated prompt optimization methods (e.g., do humans behave similar to some gradient-based optimization approach in the prompt embedding space?). There are also potential uses for crafting better image similarity metrics using the human ratings we collected.
Finally, as discussed in Section 5, our dataset and methodology is useful for assessing vision-language models. In particular, we are able to capture the ability of models to utilize feedback to adapt to a given user or tool.
@@ Appendix A Appendix
A.1 Dataset Limitations
A potential limitation in our dataset is the diversity of unique images. However, as shown in our analysis (Sections 3, 4.2), we have multiple illuminating findings despite having <200 unique images in our dataset. In collecting this dataset, there was a tradeoff – given a fixed budget to collect data, we could choose to collect more unique images or collect more user interactions per image. We opted to collect data on a fewer number of unique images with more users interacting with each image, as we believed this data would contain more insights into human interaction with the AI.
A.2 Information on Wikipedia pages scraped
@@
A.3 Examples of images generated for target parameter selection
Here we provide a few examples of images generated during target parameter selection (see Section 2.1). In Figure 9, we show the original (real) photograph (this is the target image shown to users), some examples of generated images using the caption as a prompt (i.e., the hidden goal for users), and the generated image using the selected parameters with the caption as a prompt (i.e., the generated image when a user finds the “best” prompt using the selected seed, $s_{i^{*}}$, from Equation 1). In Figure 10, we show the target image shown to users ($t_{i^{*}}$ from Equation 2), examples of generated images using the caption on various random seeds, as well as the image generated when using the seed provided to users (i.e., using the seed $s_{i_{2}^{*}}$). For both Figures, scores are normalized with respect to the rightmost image in each row – this image is guaranteed to score $100$ / $100$. Notice that the images generated from other seeds (the center two images) are also similar in quality to the image generated from the selected seed (the rightmost image).
Figure 9: Wikipedia image examples generated using the target prompt (for target selection). From left to right: (1) original Wikipedia captions (the target prompt), (2) original photograph, (3-4) images generated with random seeds, (5) image generated using the fixed seed provided to users.
Figure 10: AI-Generated image examples on the target prompt (for target selection). From left to right: (1) target prompt, (2) target image (uses a different seed than the user’s seed), (3-4) images generated with random seeds, (5) image generated using the fixed seed provided to users.
A.4 Scoring function details
In Section 2.2, we defined the scoring function, $score(x,t)$ as
$score(x,t)=\max(0,\min(100,\alpha\cdot\left|\left|\frac{C(x)}{||C(x)||_{2}}-\frac{C(t)}{||C(t)||_{2}}\right|\right|_{2}+\beta)),$
@@ A.4 Scoring function details
Since our score range is $[0,100]$, we scale the model parameters by $100$, resulting in
@@
We also add a “score adjustment” term that attempts to normalize image difficulty. In particular, for each target image, we compute the un-clipped score for the target image $t_{k}$ and the image generated using the target prompt with, $x_{k}$:
$unclipped\_score(x_{k},t_{k})=\alpha^{\prime\prime}\cdot\left|\left|\frac{C(x_{k})}{||C(x_{k})||_{2}}-\frac{C(t_{k})}{||C(t_{k})||_{2}}\right|\right|_{2}+\beta^{\prime\prime}$
This score assess the score a user would receive if they exactly entered the target prompt. We fix this value as $100$ (i.e., a perfect score prior to clipping), and set the score adjustment parameter, $c_{k}$, to appropriately normalize this score. In particular,
$c_{k}=\frac{100}{unclipped\_score(x_{k},t_{k})},$
@@
$\alpha_{k}$ $=-150.3\cdot c_{k}$
$\beta_{k}$ $=179.1\cdot c_{k}$
Note that this means each target image may have slightly different parameters as the $c_{k}$ values vary across images.
A.5 Additional information on running the game
Why we limit user input?
We deliberately limited user input to only a prompt, as opposed to giving users access to the random seed or other model hyperparameters. This was done for a few reasons:
  1. 1.
    We wanted all users who enter the same prompt for a given image to see the same output.
  2. 2.
    We wanted to limit the complexity of the task for users less/unfamiliar with text-to-image models.
  3. 3.
    We wanted users to generate new prompts and not just resample new seeds until getting lucky. While users could still employ a version of this “random resampling” strategy by making small changes to their prompts, we did not want to encourage this practice through a seed parameter. This random resampling strategy, while useful in practice, is not such an interesting result for research purposes as it is easy to simulate random resampling strategies without any user input.
Additional Technical Details
For the generative model, we use SD v2.1 [38] with the DPM Multi-step Scheduler [24] and run the model for 20 iterations. AI-generated target images use the same parameters but run for 50 iterations. 20 iterations was selected to limit latency for players to 1-3 seconds depending on the player’s internet connection. All images are generated at size 512 $\times$ 512.
@@ A.5 Additional information on running the game
A.6 Additional Stats
In Figure 12, we plot the distribution of user scores across target images. The left plot shows the initial and final scores. The shift in score to the right indicates the improvement in generated image similarity to target image. The average initial score is 56.7 (median is 57.0), and the average final score is 73.0 (median is 75.0). On the left, we plot the distribution of score improvement (the difference between the final and initial scores). The large density of 0 or close to 0 improvement is due to the “easiest” target images that users were able to generated similar examples of on their initial attempt (for example, see the first two rows of Figure 19). Most interactions result in a score difference of at most 50 points, as most users score above 50 points on their initial attempt, limiting the amount of improvement possible (the median initial score is 57.0).
In Figure 13, we plot how often a user’s score decreases, increases, or remains constant between prompts. Each point represents a series of user interactions to a generate a single target image. We note that the ratio of score increases to score decreases is similar, with the number of score increases being slightly higher. This indicates that score improvement (and so, an increase in the similarity of the generated to target image) may be effectively represented by a stochastic process with a slight bias towards score increases. This observation is the basis on which we construct our proposed steerability metric (i.e., representing score change as a stochastic process).
Figure 12: On the left: distribution of initial and final scores (across users and targets). Average initial score is 56.7; average final score is 73.0. On the right: distribution of image improvement (difference between final and initial score). Average score improvement is 16.3.
Figure 13: How score changes between queries. Each circle represents a series of user interactions to a generate a single target image. In blue, we plot the number of times a user’s score decreases over that interaction (between two queries); similarly, in orange and green we plot the number of times the user’s score increases or stays constant between prompts. We also plot best fit lines (fit using linear least squares).
A.7 ArtWhisperer-Validation Statistics
In Table 3, we provide general statistics about the ArtWhisperer-Validation Dataset. This information mirrors that statistics provided in Table 1 for the ArtWhisperer-Validation dataset. We also provide demographic information collected through Prolific in Figure 14. Users were sampled through Prolific to guarantee an even split of male and female users.
Table 3: ArtWhisperer-Validation Dataset Overview. Each row contains summary data for a different subset of the dataset. Subsets may overlap.
# Players # Target Images # Interactions Average # Prompts Average Score Median Duration Category
140 51 4572 8.14 54.77 24 s Total
73 9 762 7.47 47.55 20 s Famous person?
134 42 3937 8.19 54.58 23 s Manmade?
124 26 2154 7.69 54.61 24 s Real image?
123 24 2424 9.32 50.96 24 s Art?
94 13 1044 7.05 57.93 25 s Famous landmark?
72 9 858 8.94 50.33 26 s Nature?
70 8 730 8.11 60.25 25 s Sci-fi or space?
94 16 1501 8.25 54.21 24 s People?
123 25 2418 8.57 54.90 24 s AI image?
75 9 786 7.42 59.11 23 s City?
75 9 808 8.16 57.56 26 s Fantasy?
Figure 14: Demographic information for the paid crowd workers in the ArtWhisperer-Validation dataset.
A.8 Results for Alternative Text Embeddings
Figure 15: Same plot as Figure 6 but using the BERT [11] text embedding instead of CLIP.
In Figure 15, we replicate the analysis from Figure 6 but using a BERT [11] embedding rather than a CLIP embedding. We note similar findings and significance to the CLIP embedding analysis. We also repeated this analysis using the GLOVE [31] text embedding and again see similar results. This indicates that our analysis is robust across text representations.
A.9 Additional example images
We provide additional examples of image trajectories and diverse images in Figures 16 and 17.
@@ A.9 Additional example images
Figure 17: More examples of diverse, high-scoring prompts, as in Figure 4.
A.10 Algorithm for steerability
@@
1:  define function EstimateSteerability
2:      Input: set of images, $targetImages$
3:      Initialize array $steerability$ to track stopping times
4:      Initialize $scoreBins\leftarrow[[0,20],[21,40],[41,60],[61,80],[81,100]]$
5:      Initialize $\epsilon\leftarrow 1$
6:      for ${image}_{i}\in\text{targetImages}$ do
7:         ${markov}_{i}$leftarrow EstimateMarkovChain(${image}_{i}$, $scoreBins$, $\epsilon$)
8:         $steerability[i]$leftarrow RunMonteCarloEstimation(${markov}_{i}$)
9:      endfor
10:      Return: $\mathbf{E}[steerability]$
1:  define function EstimateMarkovChain
2:      Input: targetImage, bins, $\epsilon$
3:      Initialize bin pair count as $counts[(\text{bin}_{i},\text{bin}_{j})]\leftarrow\epsilon$, using transitions from a dummy node to model the first prompt submitted
4:      for $\text{user}_{i}\leftarrow\text{user}_{1}$ to $\text{user}_{n}$ do
5:         for $\text{score}_{i,j}\leftarrow\text{score}_{i,1}$ to $\text{score}_{i,d}$ do
6:            Convert $\text{score}_{i,j}$ to bin number, $\text{bin}_{i,j}$
7:            Increment $counts[(\text{bin}_{i-1,j},\text{bin}_{i,j})]$ by $1$
8:         endfor
9:      endfor
10:      Normalize empirical transition counts to find empirical node transition probabilities
11:      Define $markov_{\text{targetImage}}$ using the node transition probabilities
12:      Return: $markov_{\text{targetImage}}$
1:  define function RunMonteCarloEstimation
2:      Input: $markovChain$
3:      Run Monte Carlo simulation to estimate time to reach the last bin for markovChain, starting from the dummy (initial) node
4:      Return: Estimated stopping time
Algorithm 1 Image Steerability Estimation
A.11 Steerability across models